OpenStudy (anonymous):

A boardwalk is 1/4 of a rectangle which has length of 20m and a width of 12m. 1) Formulate an algebraic expression for the total area of the boardwalk given that the width of the boardwalk is unknown. 2) Form an algebraic equation with only one unknown for the total area of the boardwalk. 3) Express the equation in question 2 in the form of (x-a)(x-b)=0

OpenStudy (michele_laino):

hint: what is the area of the rectangle?

OpenStudy (anonymous):

The area of the rectangle is 240m.

OpenStudy (michele_laino):

ok! Now the area of board walk is: 240:4=...?

OpenStudy (anonymous):

60 ??

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

But how do we formulate an algebraic equation ?

OpenStudy (michele_laino):

please wait a moment, I'm checking my formulas...

OpenStudy (anonymous):

Sure.

OpenStudy (michele_laino):

for question 1), we can assume that the length of our boardwalk is 20 meters, so the width is unknown. Now I call with W the width of our boardwalk, then the product W*20 has to be equal to?

OpenStudy (anonymous):

I have my answer as 1/4(20W) = 60m

OpenStudy (michele_laino):

from your equation I get: 20*W = 240, is it right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

The total area of the boardwalk is known to be 1/4 of the area of the rectangle.

OpenStudy (michele_laino):

yes! nevertheless the area of our boardwalk has to be 60 m^2

OpenStudy (anonymous):

Yup ^^

OpenStudy (michele_laino):

for question 2) we have to find 2 numbers such that their sum is equal to the halfperimeter of our rectangle, namely (20+12)*2/2= 32, and their product has to be equal to (20*12)/4 = 60. What is that equation?

OpenStudy (michele_laino):

since we have to write a quadratic equation

OpenStudy (michele_laino):

furthermore a rectangle whose sides are: Length = 20/2 = 10 meters and width = 12/2 = 6 meters has its area equal to 60 m^2, namely equal to 240/4

OpenStudy (michele_laino):

do you know how to write that equation?

OpenStudy (anonymous):

Nope. Maybe I will just have to ask my school teacher tomorrow. Thanks for your effort :)

OpenStudy (michele_laino):

thanks! :)